# Beyond Fact Retrieval: Episodic Memory for RAG with Generative Semantic Workspaces

### v1

@@ A Probabilistic Model for Semantic Memory: The Operator Framework
GraphRAG (Edge2025GraphRAG) $\underline{0.950}\pm 0.016$ $0.657\pm 0.029$ $0.677\pm 0.034$ $0.753\pm 0.028$ $0.816\pm 0.035$ $0.781\pm 0.013$
HippoRAG2 (JimenezGutierrez2025HippoRAG) $0.829\pm 0.027$ $0.704\pm 0.029$ $\mathbf{0.817}\pm 0.026$ $\underline{0.839}\pm 0.026$ $\mathbf{0.940}\pm 0.020$ $0.812\pm 0.013$
LightRAG (Guo2024LightRAG) $0.946\pm 0.017$ $0.668\pm 0.029$ $0.615\pm 0.036$ $0.695\pm 0.031$ $0.822\pm 0.037$ $0.763\pm 0.014$
@@
GraphRAG (Edge2025GraphRAG) $\underline{0.950}\pm 0.016$ $0.764\pm 0.031$ $0.686\pm 0.035$ $0.645\pm 0.026$ $0.537\pm 0.030$ $0.748\pm 0.014$
HippoRAG2 (JimenezGutierrez2025HippoRAG) $0.829\pm 0.027$ $0.823\pm 0.026$ $\underline{0.800}\pm 0.029$ $\underline{0.749}\pm 0.026$ $\underline{0.675}\pm 0.030$ $0.787\pm 0.013$
LightRAG (Guo2024LightRAG) $0.946\pm 0.017$ $0.716\pm 0.033$ $0.628\pm 0.035$ $0.559\pm 0.029$ $0.458\pm 0.029$ $0.699\pm 0.015$
@@
GraphRAG (Edge2025GraphRAG) $\underline{0.950}\pm 0.016$ $0.625\pm 0.029$ $0.625\pm 0.034$ $0.657\pm 0.026$ $0.607\pm 0.032$ $0.714\pm 0.013$
HippoRAG2 (JimenezGutierrez2025HippoRAG) $0.829\pm 0.028$ $0.676\pm 0.028$ $\underline{0.762}\pm 0.028$ $\underline{0.754}\pm 0.025$ $\underline{0.746}\pm 0.027$ $0.753\pm 0.013$
LightRAG (Guo2024LightRAG) $0.946\pm 0.017$ $0.594\pm 0.030$ $0.587\pm 0.032$ $0.579\pm 0.028$ $0.561\pm 0.030$ $0.678\pm 0.014$
@@
GraphRAG (Edge2025GraphRAG) $\sim$7,340 $\sim$\$0.0184
HippoRAG2 (JimenezGutierrez2025HippoRAG) $\sim$8,771 $\sim$\$0.0219
LightRAG (Guo2024LightRAG) $\sim$40,476 $\sim$\$0.1012
@@
Results and Discussion
QA Performance: Table 2 presents a comparative analysis of GSW against the baseline methods detailed in Section Baseline Methods across Precision (P), Recall (R), and F1-Score (F1) metrics, categorized by the number of matching cues per query. Across the aggregated metrics, GSW achieves the highest overall F1-Score (0.850), Precision (0.865), and Recall (0.894), improving overall metrics by more than 10% over the next-best method. GSW also demonstrates consistent performance across the various Cue categories, achieving the highest score in 16 out of 18 individual metric computations, and ranking second in the remaining two, highlighting its robust performance across varying levels of episodic recall complexity. Particularly noteworthy is GSW’s performance in the ‘6+ Cues’ category. This is the most demanding scenario, where correct responses can require reasoning across information spanning up to 17 distinct chapters (see Table 1). Even in this complex setting, GSW demonstrates robust efficacy and achieves the highest performance over all metrics: F1:0.834 P:0.891, R:0.822. In particular when compared to HippoRAG2, next most performant in this category, GSW outperforms it by approximately 20% in recall. Recall, in particular, measures a framework’s ability to map queries to the correct chapter and context, and it is revealing that for all competing frameworks recall decreases as the number of matching cues increases, whereas the GSW maintains consistently strong performance, highlighting the strength of its structured representation in storing episodic information. Finally, the Vanilla LLM is consistently the poorest performing baseline (e.g overall F1 Score of 0.642) reaffirming the inherent difficulty of the episodic QA task and the necessity of specialized memory frameworks like the GSW.
Scalability on EpBench-2000: To assess the scalability of our method, we evaluate GSW on the EpBench-2000 dataset, which increases the corpus size by 10 fold. The results, presented in Table 4, show that GSW maintains its performance lead by achieving an overall F1-score of 0.773, which is 15% higher than the strongest baseline (embedding RAG), and 22% higher than other structured RAG methods. Thus, GSW’s advantages in recall and reasoning persist even at a significantly larger scale. Due to space constraints, the full breakdown table by cue category is provided in Appendix F.
@@
Table 4: Overall performance on Epbench-2000 (2000-Chapters Book). The same convention as in Table 2 is followed. For a more descriptive full table please refer to Appendix F.
@@
Token Efficiency: Beyond query performance, GSW demonstrates substantial improvements in token efficiency, as detailed in Table 3, which presents the average number of context tokens supplied to the LLM per query, and the corresponding cost for all compared methods. GSW achieves a remarkable 51% reduction in token usage when compared to the next most token-efficient baseline (GraphRAG). This advantage is even more pronounced when compared to stronger performing baselines such as Embedding RAG and HippoRAG2, against which GSW offers a token reduction of nearly 59%. GSW’s efficient approach to query resolution contributes to the reduction in token count: Rather than passing entire chapters or raw document chunks, GSW utilizes its semantic structure to generate entity-specific summaries (Prompt in Appendix A), thereby providing only targeted query-specific information to the LLM as illustrated in Appendix C. This focused contextual information also reduces hallucinations as supported by the GSW’s leading performance in the ‘0 Cues’ category, where no matching cues are present in the source document.
@@
Several additional ablation studies are presented in Appendix F, including the removal of identifier types (e.g., temporal and spatial tags), evaluations on a shortened version of the EpBench dataset, and comparisons across different retrieval strategies. These experiments highlight the contribution of each component in the GSW architecture and underscore the importance of principled memory querying. For qualitative insights into GSW’s behavior and outputs, see Appendix D.
Related Work
@@
The relevant literature has been discussed in the Introduction, and a detailed literature review is included in Appendix G. To summarize, Retrieval-Augmented Generation (RAG) (lewis_retrieval-augmented_2021; gao_retrieval-augmented_2024; karpukhin-etal-2020-dense) retrieves relevant chunks from indexed documents using dense (bert; reimers_sentence-bert_2019; lee_nv-embed_2025), sparse (robertson_probabilistic_2009), or hybrid (cormack2009reciprocal) embeddings. While effective for fact-based QA, standard RAG struggles to connect dispersed information due to its reliance on chunk-based retrieval (chen_walking_2023; merola2025reconstructing). Structured approaches like GraphRAG(Edge2025GraphRAG), LightRAG(Guo2024LightRAG) and HippoRAG(gutierrez_hipporag_2025; JimenezGutierrez2025HippoRAG) mitigate this by modeling relationships and supporting multi-hop reasoning.
Concluding Remarks and Limitations
@@
Nevertheless, we identify key limitations and avenues for future work. Firstly, GSW’s evaluation, while utilizing EpBench for its strengths in spatiotemporal assessment, is constrained by the limited scope of current episodic memory benchmarks in thoroughly probing the complex evolution of actor roles and states within extended narratives; we are developing a more comprehensive benchmark to specifically address this gap. Secondly, the present GSW implementation relies on a strong closed-source LLM (GPT-4o). Empirical validation of promising open-source alternatives (yang2024qwen2; grattafiori2024llama; team2025gemma) within our Operator-Reconciler architecture is therefore essential. Expanding GSW to process and integrate information from diverse data modalities beyond text also presents an important direction for future development, aiming to broaden its applicability to more complex, real-world experiential inputs.
@@ Appendix C GSW QA Example
Appendix D Qualitative Analysis of GSW performance
@@ Appendix E Implementation Details
For HippoRAG2 (JimenezGutierrez2025HippoRAG), GraphRAG (Edge2025GraphRAG), and LightRAG (Guo2024LightRAG), we adhere to each method’s default hyperparameters and prompt formats as provided in their respective official implementations. To ensure consistency across baselines, we modify the answering model in HippoRAG2 to use GPT-4o, aligning it with other evaluated methods. Additionally, we set top-k to 17 for HippoRAG2 to retrieve the top 17 relevant documents to align with the QA settings. The detailed configurations for each baseline are listed in Tables79.
@@ Appendix F Ablation Studies
GraphRAG (Edge2025GraphRAG) $\mathbf{0.943}\pm 0.025$ $0.747\pm 0.038$ $0.639\pm 0.040$ $0.692\pm 0.038$ $0.795\pm 0.043$ $0.761\pm 0.017$
HippoRAG2 (JimenezGutierrez2025HippoRAG) $0.620\pm 0.051$ $0.638\pm 0.032$ $0.803\pm 0.032$ $0.824\pm 0.028$ $\underline{0.893}\pm 0.021$ $0.759\pm 0.016$
LightRAG (Guo2024LightRAG) $0.790\pm 0.042$ $0.534\pm 0.039$ $0.560\pm 0.040$ $0.593\pm 0.035$ $0.787\pm 0.039$ $0.649\pm 0.018$
@@
GraphRAG (Edge2025GraphRAG) $\mathbf{0.943}\pm 0.025$ $0.492\pm 0.037$ $0.587\pm 0.039$ $0.538\pm 0.036$ $0.321\pm 0.025$ $0.548\pm 0.017$
HippoRAG2 (JimenezGutierrez2025HippoRAG) $0.620\pm 0.050$ $0.703\pm 0.034$ $0.769\pm 0.031$ $\underline{0.647}\pm 0.029$ $\underline{0.491}\pm 0.026$ $0.648\pm 0.016$
LightRAG (Guo2024LightRAG) $0.790\pm 0.042$ $0.525\pm 0.038$ $0.549\pm 0.038$ $0.440\pm 0.033$ $0.270\pm 0.017$ $0.497\pm 0.017$
@@
GraphRAG (Edge2025GraphRAG) $\mathbf{0.943}\pm 0.025$ $0.436\pm 0.035$ $0.547\pm 0.038$ $0.541\pm 0.036$ $0.405\pm 0.027$ $0.544\pm 0.017$
HippoRAG2 (JimenezGutierrez2025HippoRAG) $0.620\pm 0.050$ $0.583\pm 0.031$ $0.732\pm 0.031$ $\underline{0.681}\pm 0.027$ $\underline{0.578}\pm 0.026$ $0.635\pm 0.015$
LightRAG (Guo2024LightRAG) $0.790\pm 0.042$ $0.436\pm 0.034$ $0.514\pm 0.037$ $0.463\pm 0.033$ $0.375\pm 0.021$ $0.494\pm 0.016$
@@
Appendix G Related work on Memory Augmentation for LLMs
Enabling LLMs to effectively process long narratives requires capabilities akin to human episodic memory – constructing and maintaining a dynamic, coherent understanding of events unfolding over space and time (tulving_episodic_1972; eichenbaum_corticalhippocampal_2000). Key to this is the ability to accurately track entities, including their evolving states and roles, and to ground events and answer queries based on specific spatial and temporal contexts established within the narrative (huet_episodic_2025). While LLMs possess remarkable core abilities, achieving this level of sophisticated, stateful reasoning over extended sequences remains a significant challenge. The following sections analyze inherent limitations in common approaches used to provide context to LLMs, evaluating why they often fall short of systematically delivering these specific episodic memory capabilities.
@@ Appendix G Related work on Memory Augmentation for LLMs
One approach to providing LLMs with relevant context is to leverage their increasingly large context windows, potentially feeding the entire long narrative along with a query into the prompt. The rapid expansion of context lengths, now reaching millions of tokens, has certainly broadened the scope of tasks LLMs can handle by allowing more raw information to be processed simultaneously (team2024gemini).
However, relying solely on this native processing mechanism faces significant hurdles when evaluated against the demands of episodic memory. Firstly, while context windows are growing, they are not infinite, and extremely long narratives may still exceed even the largest available limits. Secondly, even when a narrative technically fits, processing vast amounts of text for every query is computationally expensive, impacting latency and cost. More fundamentally, processing quality often degrades with extreme context lengths (leng2024long; hsieh_ruler_2024; wang2024multimodal). Research has shown that LLMs can struggle to consistently access and utilize information spread across very long contexts, with performance notably dipping for information located in the middle (lost in the middle phenomenon) (liu_lost_2023). Feeding potentially large amounts of irrelevant text alongside the crucial details for a specific episodic query can distract the model and hinder its ability to pinpoint and reason over the necessary information.
Finally, perhaps the most critical limitation for systematic episodic tracking is the inherently unstructured nature of the input context. Even with all the necessary details about entity states, roles, locations, and times present within the text, the LLM lacks explicit mechanisms to structure this information dynamically. It must rely solely on its attention mechanism and in-context learning to piece together relationships, track state changes, and maintain temporal coherence across potentially thousands of tokens. This makes the reliable, systematic tracking required for robust episodic memory challenging and often brittle when relying only on the native context window (huet_episodic_2025).
@@
To overcome the challenges of static parametric knowledge and the inefficiencies of processing entire documents in context, Retrieval-Augmented Generation (RAG) has become a standard technique (lewis_retrieval-augmented_2021; gao_retrieval-augmented_2024). The typical RAG pipeline involves pre-processing a knowledge corpus (e.g., the entire narrative document) into smaller chunks. These chunks are then indexed, commonly using dense vector embeddings obtained from encoder style LLMs(bert; reimers_sentence-bert_2019; lee_nv-embed_2025), though sparse methods like BM25(robertson_probabilistic_2009) or hybrid approaches are also employed (cormack2009reciprocal). At inference time, the user query is used to retrieve the top-k most relevant chunks from the index based on a similarity metric (e.g., cosine similarity for dense vectors). These retrieved chunks are then presented as augmented context to an LLM, which generates the final response based on both its parametric knowledge and the retrieved information.(ram_-context_2023)
This approach has proven effective for many knowledge-intensive tasks, particularly fact-based question answering where retrieving specific evidence snippets is sufficient (karpukhin-etal-2020-dense). However, when evaluated against the requirements of robust episodic memory recall over long narratives, the limitations of standard RAG become apparent (huet_episodic_2025). Firstly, the process of retrieving discrete, potentially disconnected chunks based on local query relevance often fragments the narrative flow. This makes it exceedingly difficult for the LLM to reliably follow evolving storylines or track the changing states and roles of entities over time, as the necessary context may be spread across multiple chunks that are not retrieved together(chen_walking_2023).
Moreover, this fragmentation problem is compounded by the framework being highly sensitive to the initial chunking strategy(merola2025reconstructing). Arbitrary chunk boundaries can split crucial information related to an event or an entity’s state, leading to information loss during retrieval. For instance, if a character’s state changes within a passage, but the chunking algorithm divides this passage at an inopportune point, the complete context of this state change may not be captured in any single retrieved chunk. Optimal chunking is non-trivial and can significantly impact the ability to reconstruct the necessary context for complex episodic reasoning. Consequently, while standard RAG offers efficiency gains over naive long-context processing, its inherent lack of structure and narrative coherence makes it ill-suited for systematically addressing the dynamic, stateful, and context-dependent nature of episodic memory tasks.

### v2

@@ 3.4 Implementation Details
GraphRAG (Edge et al. 2025) $\underline{0.950}\pm 0.016$ $0.657\pm 0.029$ $0.677\pm 0.034$ $0.753\pm 0.028$ $0.816\pm 0.035$ $0.781\pm 0.013$
HippoRAG2 (Gutiérrez et al. 2025b) $0.829\pm 0.027$ $0.704\pm 0.029$ $\mathbf{0.817}\pm 0.026$ $\underline{0.839}\pm 0.026$ $\mathbf{0.940}\pm 0.020$ $0.812\pm 0.013$
LightRAG (Guo et al. 2025) $0.946\pm 0.017$ $0.668\pm 0.029$ $0.615\pm 0.036$ $0.695\pm 0.031$ $0.822\pm 0.037$ $0.763\pm 0.014$
@@
GraphRAG (Edge et al. 2025) $\underline{0.950}\pm 0.016$ $0.764\pm 0.031$ $0.686\pm 0.035$ $0.645\pm 0.026$ $0.537\pm 0.030$ $0.748\pm 0.014$
HippoRAG2 (Gutiérrez et al. 2025b) $0.829\pm 0.027$ $0.823\pm 0.026$ $\underline{0.800}\pm 0.029$ $\underline{0.749}\pm 0.026$ $\underline{0.675}\pm 0.030$ $0.787\pm 0.013$
LightRAG (Guo et al. 2025) $0.946\pm 0.017$ $0.716\pm 0.033$ $0.628\pm 0.035$ $0.559\pm 0.029$ $0.458\pm 0.029$ $0.699\pm 0.015$
@@
GraphRAG (Edge et al. 2025) $\underline{0.950}\pm 0.016$ $0.625\pm 0.029$ $0.625\pm 0.034$ $0.657\pm 0.026$ $0.607\pm 0.032$ $0.714\pm 0.013$
HippoRAG2 (Gutiérrez et al. 2025b) $0.829\pm 0.028$ $0.676\pm 0.028$ $\underline{0.762}\pm 0.028$ $\underline{0.754}\pm 0.025$ $\underline{0.746}\pm 0.027$ $0.753\pm 0.013$
LightRAG (Guo et al. 2025) $0.946\pm 0.017$ $0.594\pm 0.030$ $0.587\pm 0.032$ $0.579\pm 0.028$ $0.561\pm 0.030$ $0.678\pm 0.014$
@@
GraphRAG (Edge et al. 2025) $\sim$7,340 $\sim$\$0.0184
HippoRAG2 (Gutiérrez et al. 2025b) $\sim$8,771 $\sim$\$0.0219
LightRAG (Guo et al. 2025) $\sim$40,476 $\sim$\$0.1012
@@
4 Results and Discussion
QA Performance: Table 2 presents a comparative analysis of GSW against the baseline methods detailed in Section 3.3 across Precision (P), Recall (R), and F1-Score (F1) metrics, categorized by the number of matching cues per query. Across the aggregated metrics, GSW achieves the highest overall F1-Score (0.850), Precision (0.865), and Recall (0.894), improving overall metrics by more than 10% over the next-best method. GSW also demonstrates consistent performance across the various Cue categories, achieving the highest score in 16 out of 18 individual metric computations, and ranking second in the remaining two, highlighting its robust performance across varying levels of episodic recall complexity. Particularly noteworthy is GSW’s performance in the ‘6+ Cues’ category. This is the most demanding scenario, where correct responses can require reasoning across information spanning up to 17 distinct chapters (see Table 1). Even in this complex setting, GSW demonstrates robust efficacy and achieves the highest performance over all metrics: F1:0.834 P:0.891, R:0.822. In particular when compared to HippoRAG2, next most performant in this category, GSW outperforms it by approximately 20% in recall. Recall, in particular, measures a framework’s ability to map queries to the correct chapter and context, and it is revealing that for all competing frameworks recall decreases as the number of matching cues increases, whereas the GSW maintains consistently strong performance, highlighting the strength of its structured representation in storing episodic information. Finally, the Vanilla LLM is consistently the poorest performing baseline (e.g overall F1 Score of 0.642) reaffirming the inherent difficulty of the episodic QA task and the necessity of specialized memory frameworks like the GSW.
Scalability on EpBench-2000: To assess the scalability of our method, we evaluate GSW on the EpBench-2000 dataset, which increases the corpus size by 10 fold. The results, presented in Table 4, show that GSW maintains its performance lead by achieving an overall F1-score of 0.773, which is 15% higher than the strongest baseline (embedding RAG), and 22% higher than other structured RAG methods. Thus, GSW’s advantages in recall and reasoning persist even at a significantly larger scale.
@@ 4 Results and Discussion
Table 4: Overall performance on Epbench-2000 (2000-Chapters Book). The same convention as in Table 2 is followed.
@@
Token Efficiency: Beyond query performance, GSW demonstrates substantial improvements in token efficiency, as detailed in Table 3, which presents the average number of context tokens supplied to the LLM per query, and the corresponding cost for all compared methods. GSW achieves a remarkable 51% reduction in token usage when compared to the next most token-efficient baseline (GraphRAG). This advantage is even more pronounced when compared to stronger performing baselines such as Embedding RAG and HippoRAG2, against which GSW offers a token reduction of nearly 59%. GSW’s efficient approach to query resolution contributes to the reduction in token count: Rather than passing entire chapters or raw document chunks, GSW utilizes its semantic structure to generate entity-specific summaries, thereby providing only targeted query-specific information to the LLM. This focused contextual information also reduces hallucinations as supported by the GSW’s leading performance in the ‘0 Cues’ category, where no matching cues are present in the source document.
@@
Several additional ablation studies are presented in the full version, including the removal of identifier types (e.g., temporal and spatial tags), evaluations on a shortened version of the EpBench dataset, and comparisons across different retrieval strategies. These experiments highlight the contribution of each component in the GSW architecture and underscore the importance of principled memory querying. For qualitative insights into GSW’s behavior and outputs, see the full version.
5 Related Work
The relevant literature has been discussed in the Introduction, and a detailed literature review is included in the full version. While modern LLMs offer increasingly large context windows, processing quality degrades with extreme lengths (Leng et al. 2024; Hsieh et al. 2024), with performance notably dipping for information in the middle of long contexts (Liu et al. 2023). This makes reliable episodic tracking challenging when relying solely on native context windows.
@@ 5 Related Work
Retrieval-Augmented Generation (RAG) (Lewis et al. 2021; Gao et al. 2024; Karpukhin et al. 2020) addresses this by retrieving relevant chunks using dense (Devlin et al. 2019; Reimers and Gurevych 2019; Lee et al. 2025), sparse (Robertson and Zaragoza 2009), or hybrid (Cormack et al. 2009) embeddings. While effective for fact-based QA, standard RAG struggles to connect dispersed information due to chunk-based retrieval (Chen et al. 2023; Merola and Singh 2025). Structured approaches like GraphRAG (Edge et al. 2025), LightRAG (Guo et al. 2025) and HippoRAG (Gutiérrez et al. 2025a, b) mitigate this by modeling relationships and supporting multi-hop reasoning.
Other research efforts have targeted episodic memory more directly. Larimar (Das et al. 2024) proposes modifications to the LLM’s attention mechanism, while EM-LLM (Fountas et al. 2024) introduces memory components integrated with open-weight models. While promising, these approaches often require architectural changes or are designed for specific models. In contrast, GSW is a plug-and-play module compatible with any LLM, requiring no specialized training or fine-tuning.
6 Concluding Remarks and Limitations
@@
Nevertheless, we identify key limitations and avenues for future work. Firstly, GSW’s evaluation, while utilizing EpBench for its strengths in spatiotemporal assessment, is constrained by the limited scope of current episodic memory benchmarks in thoroughly probing the complex evolution of actor roles and states within extended narratives; we are developing a more comprehensive benchmark to address this gap. Secondly, the present GSW implementation relies on a strong closed-source LLM (GPT-4o). Empirical validation of promising open-source alternatives (Yang et al. 2024; Grattafiori et al. 2024) within our framework is essential. Expanding GSW to diverse data modalities beyond text is also an important direction for future work.
@@ C GSW QA Example
D Qualitative Analysis of GSW performance
@@ E Implementation Details
For HippoRAG2 (Gutiérrez et al. 2025b), GraphRAG (Edge et al. 2025), and LightRAG (Guo et al. 2025), we adhere to each method’s default hyperparameters and prompt formats as provided in their respective official implementations. To ensure consistency across baselines, we modify the answering model in HippoRAG2 to use GPT-4o, aligning it with other evaluated methods. Additionally, we set top-k to 17 for HippoRAG2 to retrieve the top 17 relevant documents to align with the QA settings. The detailed configurations for each baseline are listed in Tables79.
@@ F Ablation Studies
GraphRAG (Edge et al. 2025) $\mathbf{0.943}\pm 0.025$ $0.747\pm 0.038$ $0.639\pm 0.040$ $0.692\pm 0.038$ $0.795\pm 0.043$ $0.761\pm 0.017$
HippoRAG2 (Gutiérrez et al. 2025b) $0.620\pm 0.051$ $0.638\pm 0.032$ $0.803\pm 0.032$ $0.824\pm 0.028$ $\underline{0.893}\pm 0.021$ $0.759\pm 0.016$
LightRAG (Guo et al. 2025) $0.790\pm 0.042$ $0.534\pm 0.039$ $0.560\pm 0.040$ $0.593\pm 0.035$ $0.787\pm 0.039$ $0.649\pm 0.018$
@@
GraphRAG (Edge et al. 2025) $\mathbf{0.943}\pm 0.025$ $0.492\pm 0.037$ $0.587\pm 0.039$ $0.538\pm 0.036$ $0.321\pm 0.025$ $0.548\pm 0.017$
HippoRAG2 (Gutiérrez et al. 2025b) $0.620\pm 0.050$ $0.703\pm 0.034$ $0.769\pm 0.031$ $\underline{0.647}\pm 0.029$ $\underline{0.491}\pm 0.026$ $0.648\pm 0.016$
LightRAG (Guo et al. 2025) $0.790\pm 0.042$ $0.525\pm 0.038$ $0.549\pm 0.038$ $0.440\pm 0.033$ $0.270\pm 0.017$ $0.497\pm 0.017$
@@
GraphRAG (Edge et al. 2025) $\mathbf{0.943}\pm 0.025$ $0.436\pm 0.035$ $0.547\pm 0.038$ $0.541\pm 0.036$ $0.405\pm 0.027$ $0.544\pm 0.017$
HippoRAG2 (Gutiérrez et al. 2025b) $0.620\pm 0.050$ $0.583\pm 0.031$ $0.732\pm 0.031$ $\underline{0.681}\pm 0.027$ $\underline{0.578}\pm 0.026$ $0.635\pm 0.015$
LightRAG (Guo et al. 2025) $0.790\pm 0.042$ $0.436\pm 0.034$ $0.514\pm 0.037$ $0.463\pm 0.033$ $0.375\pm 0.021$ $0.494\pm 0.016$
@@
G Related work on Memory Augmentation for LLMs
Enabling LLMs to effectively process long narratives requires capabilities akin to human episodic memory – constructing and maintaining a dynamic, coherent understanding of events unfolding over space and time (Tulving 1972; Eichenbaum 2000). Key to this is the ability to accurately track entities, including their evolving states and roles, and to ground events and answer queries based on specific spatial and temporal contexts established within the narrative (Huet et al. 2025). While LLMs possess remarkable core abilities, achieving this level of sophisticated, stateful reasoning over extended sequences remains a significant challenge. The following sections analyze inherent limitations in common approaches used to provide context to LLMs, evaluating why they often fall short of systematically delivering these specific episodic memory capabilities.
@@ G Related work on Memory Augmentation for LLMs
One approach to providing LLMs with relevant context is to leverage their increasingly large context windows, potentially feeding the entire long narrative along with a query into the prompt. The rapid expansion of context lengths, now reaching millions of tokens, has certainly broadened the scope of tasks LLMs can handle by allowing more raw information to be processed simultaneously (Team et al. 2024).
However, relying solely on this native processing mechanism faces significant hurdles when evaluated against the demands of episodic memory. Firstly, while context windows are growing, they are not infinite, and extremely long narratives may still exceed even the largest available limits. Secondly, even when a narrative technically fits, processing vast amounts of text for every query is computationally expensive, impacting latency and cost. More fundamentally, processing quality often degrades with extreme context lengths (Leng et al. 2024; Hsieh et al. 2024; Wang et al. 2024). Research has shown that LLMs can struggle to consistently access and utilize information spread across very long contexts, with performance notably dipping for information located in the middle (lost in the middle phenomenon) (Liu et al. 2023). Feeding potentially large amounts of irrelevant text alongside the crucial details for a specific episodic query can distract the model and hinder its ability to pinpoint and reason over the necessary information.
Finally, perhaps the most critical limitation for systematic episodic tracking is the inherently unstructured nature of the input context. Even with all the necessary details about entity states, roles, locations, and times present within the text, the LLM lacks explicit mechanisms to structure this information dynamically. It must rely solely on its attention mechanism and in-context learning to piece together relationships, track state changes, and maintain temporal coherence across potentially thousands of tokens. This makes the reliable, systematic tracking required for robust episodic memory challenging and often brittle when relying only on the native context window (Huet et al. 2025).
@@
To overcome the challenges of static parametric knowledge and the inefficiencies of processing entire documents in context, Retrieval-Augmented Generation (RAG) has become a standard technique (Lewis et al. 2021; Gao et al. 2024). The typical RAG pipeline involves pre-processing a knowledge corpus (e.g., the entire narrative document) into smaller chunks. These chunks are then indexed, commonly using dense vector embeddings obtained from encoder style LLMs(Devlin et al. 2019; Reimers and Gurevych 2019; Lee et al. 2025), though sparse methods like BM25(Robertson and Zaragoza 2009) or hybrid approaches are also employed (Cormack et al. 2009). At inference time, the user query is used to retrieve the top-k most relevant chunks from the index based on a similarity metric (e.g., cosine similarity for dense vectors). These retrieved chunks are then presented as augmented context to an LLM, which generates the final response based on both its parametric knowledge and the retrieved information.(Ram et al. 2023)
This approach has proven effective for many knowledge-intensive tasks, particularly fact-based question answering where retrieving specific evidence snippets is sufficient (Karpukhin et al. 2020). However, when evaluated against the requirements of robust episodic memory recall over long narratives, the limitations of standard RAG become apparent (Huet et al. 2025). Firstly, the process of retrieving discrete, potentially disconnected chunks based on local query relevance often fragments the narrative flow. This makes it exceedingly difficult for the LLM to reliably follow evolving storylines or track the changing states and roles of entities over time, as the necessary context may be spread across multiple chunks that are not retrieved together(Chen et al. 2023).
Moreover, this fragmentation problem is compounded by the framework being highly sensitive to the initial chunking strategy(Merola and Singh 2025). Arbitrary chunk boundaries can split crucial information related to an event or an entity’s state, leading to information loss during retrieval. For instance, if a character’s state changes within a passage, but the chunking algorithm divides this passage at an inopportune point, the complete context of this state change may not be captured in any single retrieved chunk. Optimal chunking is non-trivial and can significantly impact the ability to reconstruct the necessary context for complex episodic reasoning. Consequently, while standard RAG offers efficiency gains over naive long-context processing, its inherent lack of structure and narrative coherence makes it ill-suited for systematically addressing the dynamic, stateful, and context-dependent nature of episodic memory tasks.